Be glad if someone could help me with this. Thanks
If an={-3,2,7,12,17...}
Determine the sum of the firts 40 terms of an

Answer by rothauserc(4718)

this is an arithmetic sequence
the summation of the first n terms = (n/2)(2a + d(n-1)) where a is the first term and d is the difference between terms
*****************************************************************************
In our problem, n = 40, d = 5 and a = -3
summation of first 40 terms = (40/2)(2*(-3) + 5(40-1)) = 3780